# Statistical Taylor Expansion111

### v3

@@ 2.3 Statistical Taylor Expansion
An analytic function $f(x)$ can be precisely calculated in a range by the Taylor series as shown in Formula (2.5). Formula (2.4) defines variance momentum $\zeta(n)$. Using Formula (2.4), Formula (2.6) and (2.3) gives the mean $\overline{f(x)}$ and the variance $\delta^{2}f(x)$ for $f(x)$, respectively. $\overline{f(x)}-f(x)$ is defined as the uncertainty bias which is the effect of the input uncertainty to the result value.
@@
$\tilde{y}=$ $f(x+\tilde{x})=f(x+\tilde{z}\delta x)=\sum_{n=0}^{\infty}\frac{f^{(n)}_{x}}{n!}\tilde{z}^{n}(\delta x)^{n};$ (2.5)
@@
$=\sum_{n=2}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\frac{f^{(j)}_{x}}{j!}\frac{f^{(n-j)}_{x}}{(n-j)!}\big{(}\zeta(n)-\zeta(j)\zeta(n-j)\big{)};$ (2.7)
When $\sigma$ is limited, Formula (2.4) provides statistical bounding $\tilde{x}\in(x-\sigma\delta x,x+\sigma\delta x)$. The probability $\epsilon$ for $\tilde{x}\not\in(x-\sigma\delta x,x+\sigma\delta x)$ is defined as bounding leakage.
@@
$\delta^{2}f(x,y)=$ $\sum_{m=2}^{\infty}\sum_{n=2}^{\infty}(\delta x)^{m}(\delta y)^{n}\sum_{i=0}^{m}\sum_{j=0}^{n}\frac{f^{(i,j)}_{(x,y)}}{i!\;j!}\frac{f^{(m-i,n-j)}_{(x,y)}}{(m-i)!\;(n-j)!}$
@@ 2.4 Bounding Symmetry and Asymptote
A bounding $[x-\sigma_{-}\delta x,x+\sigma_{+}\delta x]$ is mean-preserving if $\zeta(1)=0$, which has the same mean as the unbounded distribution. Usually, the two positive value $\sigma_{-}$ and $\sigma_{+}$ can be chosen to achieve mean-preserving. With mean-preserving bounding, Formula (2.11) and (2.12) give the result for $x\pm y$, while Formula (2.13) and (2.14) give the result for $xy$:
@@ 2.4.3 An Input Uncertainty with Limited Range
Formula (2.20) shows a distribution density function with $\tilde{x}\in[0,+\infty)$, whose bounding range $[\sigma_{-}\lambda,\sigma\lambda]$ satisfying $0<\sigma_{-}<1<\sigma$. Formula (2.21) shows its mean-preserving equation. Formula (2.22) shows its variance momentum. Formula (2.23) shows its asymptote. Formula (2.24) shows its bounding leakage.
@@
$e^{-(\sigma_{-})}(\sigma_{-})^{2}=e^{-\sigma}\sigma^{2};$ (2.21)
$\zeta(n)$ $=(n+1)!\left(e^{-(\sigma_{-})}\sum_{j=0}^{n+1}\frac{(\sigma_{-})^{j}}{j!}-e^{-\sigma}\sum_{j=0}^{n+1}\frac{\sigma^{j}}{j!}\right);$ (2.22)
@@
$\epsilon$ $=1-e^{-(\sigma_{-})}+e^{-\sigma};$ (2.24)
@@ 2.5 One Dimensional Examples
$\frac{\delta^{2}x^{c}}{(x^{c})^{2}}$ $=\sum_{n=2}^{\infty}P(x)^{n}\sum_{j=1}^{n-1}\begin{pmatrix}c\\ j\end{pmatrix}\begin{pmatrix}c\\ n-j\end{pmatrix}\left(\zeta(n)-\zeta(j)\zeta(n-j)\right);$ (2.33)
@@
$\overline{\sin(x)}=$ $sinx(x)+\sum_{n=1}^{\infty}(\delta x)^{n}\eta(n,x)\frac{\zeta(n)}{n!};$ (2.35)
$\delta^{2}\sin(x)=$ $\sum_{n=2}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\frac{\eta(j,x)\eta(n-j,x)}{j!(n-j)!}\left(\zeta(n)-\zeta(j)\zeta(n-j)\right);$ (2.36)
@@ 3.1 Choices of Bounding Factors
For discussion simplicity, only Gaussian input uncertainty will be considered in this paper.
For Gaussian input uncertainty, when $\sigma=5$, $\epsilon=5.73\times 10^{-7}$ according to Formula (2.19), which is small enough for most applications. $\sigma=5$ is also a golden standard to assert a negative result [3]. Thus, $\sigma=5$ in variance arithmetic for Gaussian input uncertainty.
@@ 3.2 Low-Order Approximation
To keep the result precision finer than $1$ for $e^{x\pm\delta x}$, $\delta x\lesssim 0.833$ regardless of $x$. In contrast, the result upper bound uncertainty $\delta\log(x\pm\delta x)\simeq 0.212$. Both are expected from the relation $\delta x\rightarrow P(e^{x})$, and $P(x)\rightarrow\delta\log(x)$, respectively.
@@
3.3 Convergence, Stability, Validity, and Reliability
[…]
figure The measured upper bound $\delta x$ for $\log{x\pm\delta x}$ for different $x$ as shown by the x-axis, using the y-axis on the left. Also shown are the corresponding uncertainty bias and uncertainty, using the y-axis on the right.
@@ 3.3 Convergence, Stability, Validity, and Reliability
Due to digital limitation of conventional floating-point representation, $\zeta(2n)$ can only be calculated to $2442n<$ 244 before becoming infinitive, Thus, variance arithmetic has different convergence rule from its counterpart of Taylor expansion of precise variable. In variance arithmetic, an infinitive expansion (including Taylor expansion) can fail for the following reasons: not monotonic, not stable, not positive, not finite, and not reliable.
@@
According to Formula (3.3), the condition for Formula (2.30) for $\log(x\pm\delta x)$ to converge is $P(x)\leq 1/\sigma=1/5$, which is confirmed by Figure 3.3. Variance arithmetic rejects the distributional zero of $\log(0)$ inside the range of $[x-\delta x,x+\delta x]$ statistically due to the divergence of Formula (2.30) mathematically, with $\zeta(2n)$ as the connection of these two worlds.
According to Formula (3.3), except when c is a nature number, Formula (2.33) for $(x\pm\delta x)^{c}$ converges near $P(x)\simeq 1/\sigma=1/5$, with the upper bound $P(x)$ increasing with $c$. This is confirmed by Figure 3.3. However, due to complexity of Formula (2.33), there is yet any quantitative formula between $\overline{P(x)}$ and $c$. Qualitatively, $\delta^{2}1/x$ converges slower than $\delta^{2}\sqrt{x}$ according to Formula (3.10) and (3.8).
@@
Conventional condition to terminate an infinitive series numerically is when the residual of an expansion is less than the unit in the last place of the result [13]. In practice, the last expansion term usually proxies for the expansion residual if either the residual is known to be less than the last expansion term, or if the residual is difficult to calculate [13]. This method no longer works in variance arithmetic because the expansion is limited to $2442n<$ 244, and because the expansion result is no longer precise. The probability of an imprecie value $x\pm\delta x$ not equal 00 is indicated by its z-statistic $\frac{|x|}{\delta x}$ [5]. If this probability is the bounding leakage $\epsilon$, Formula (3.13) gives the corresponding upper bound $\tau$ for this z-statistic. Because variance arithmetic has already guaranteed that the expansion is absolutely and monotonically decreasing, if the absolute value of the last expansion term is more than $\tau$-fold of the result uncertainty, the result is invalidated as not stable.
@@
Formula (2.32) shows that $\overline{x^{c}}$ converges slower with decreasing $c$. In Figure 3.3, beyond the upper bound $\delta x$, the expansion is not stable when $c\leq-1.42$, or not monotonic when $c>-1.41$.
@@
If the result of Taylor expansion is not finite, the expansion is invalidated as not finite.
Because $\eta(2n)\ll 2n!!$!, Formula (2.27) for $e^{x\pm\delta x}$ always converge until the result becomes not finite, such as $e^{0\pm 8.001}$.
@@ 3.4 Continuity
When $c$ is a nature number $n$, Formula (2.33) has $n$ terms, so that $\delta x$ has no upper bound for $(x\pm\delta x)^{n}$ unless for not finite, such as Formula (3.6) for $(x\pm\delta x)^{2}$. The result mean, variance and histogram of $(x\pm\delta x)^{c}$ are continuous around $c=n$.

### v4

@@ 2.3 Statistical Taylor Expansion
An analytic function $f(x)$ can be precisely calculated in a range by the Taylor series as shown in Formula (2.5). Formula (2.4) defines bounded momentum $\zeta(n)$, in which $0<\varrho,\sigma$ are bounding factors. Using Formula (2.4), Formula (2.6) and (2.3) gives the mean $\overline{f(x)}$ and the variance $\delta^{2}f(x)$ for $f(x)$, respectively. $\overline{f(x)}-f(x)$ is defined as the uncertainty bias which is the effect of the input uncertainty to the result value.
@@
$f(x+\tilde{x})=$ $f(x+\tilde{z}\delta x)=\sum_{n=0}^{\infty}\frac{f^{(n)}_{x}}{n!}\tilde{z}^{n}(\delta x)^{n};$ (2.5)
@@
$=\sum_{n=1}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\frac{f^{(j)}_{x}}{j!}\frac{f^{(n-j)}_{x}}{(n-j)!}\big{(}\zeta(n)-\zeta(j)\zeta(n-j)\big{)};$ (2.7)
@@
$\delta^{2}f(x,y)=$ $\sum_{m=1}^{\infty}\sum_{n=1}^{\infty}(\delta x)^{m}(\delta y)^{n}\sum_{i=0}^{m}\sum_{j=0}^{n}\frac{f^{(i,j)}_{(x,y)}}{i!\;j!}\frac{f^{(m-i,n-j)}_{(x,y)}}{(m-i)!\;(n-j)!}$
@@
2.4 Statistical Bounding
When the bounding factors $\varrho$ and $\sigma$ are limited, Formula (2.4) provides statistical bounding $\tilde{x}\in[x-\varrho\delta x,x+\sigma\delta x]$. The probability $\epsilon$ for $\tilde{x}\not\in[x-\varrho\delta x,x+\sigma\delta x]$ is defined as bounding leakage.
@@ 2.5 Bounding Symmetry and Asymptote
A bounding $[x-\varrho\delta x,x+\sigma\delta x]$ is mean-preserving if $\zeta(1)=0$, which has the same mean as the unbounded distribution. To achieve mean-preserving, $\sigma$ determines $\varrho$. With mean-preserving bounding, Formula (2.11) and (2.12) give the result for $x\pm y$, while Formula (2.13) and (2.14) give the result for $xy$:
@@ 2.5.3 An Input Uncertainty with Limited Range
Formula (2.20) shows a distribution density function with $\tilde{x}\in[0,+\infty)$, whose bounding range $[\varrho\lambda,\sigma\lambda]$ satisfying $0<\varrho<1<\sigma$. Formula (2.21) shows its mean-preserving equation. Formula (2.22) shows its bounded momentum. Formula (2.23) shows its asymptote. Formula (2.24) shows its bounding leakage.
@@
$e^{-\varrho}\varrho^{2}=e^{-\sigma}\sigma^{2};$ (2.21)
$\zeta(n)$ $=(n+1)!\left(e^{-\varrho}\sum_{j=0}^{n+1}\frac{\varrho^{j}}{j!}-e^{-\sigma}\sum_{j=0}^{n+1}\frac{\sigma^{j}}{j!}\right);$ (2.22)
@@
$\epsilon$ $=1-e^{-\varrho}+e^{-\sigma};$ (2.24)
@@ 2.6 One Dimensional Examples
$\frac{\delta^{2}x^{c}}{(x^{c})^{2}}$ $=\sum_{n=2}^{\infty}P(x)^{n}\sum_{j=1}^{n-1}\begin{pmatrix}c\\ j\end{pmatrix}\begin{pmatrix}c\\ n-j\end{pmatrix}\big{(}\zeta(n)-\zeta(j)\zeta(n-j)\big{)};$ (2.33)
@@
$\overline{\sin(x)}=$ $\sin(x)+\sum_{n=1}^{\infty}(\delta x)^{n}\eta(n,x)\frac{\zeta(n)}{n!};$ (2.35)
$\delta^{2}\sin(x)=$ $\sum_{n=2}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\frac{\eta(j,x)\eta(n-j,x)}{j!(n-j)!}\big{(}\zeta(n)-\zeta(j)\zeta(n-j)\big{)};$ (2.36)
@@ 3.1 Choices of Bounding Factors
For discussion simplicity, only Gaussian input uncertainty will be considered in this paper. For Gaussian input uncertainty, when $\sigma=5$, $\epsilon=5.73\times 10^{-7}$ according to Formula (2.19), which is small enough for most applications. $\sigma=5$ is also a golden standard to assert a negative result [3]. Thus, $\sigma=5$ in variance arithmetic for Gaussian input uncertainty.
@@ 3.2 Low-Order Approximation
3.3 Expansion Calculation
@@
Due to digital limitation of conventional floating-point representation, $\zeta(2n)$ can only be calculated to $4422n<$ 442 before becoming infinitive, Thus, variance arithmetic has different convergence rule from its counterpart of Taylor expansion of precise variable. In variance arithmetic, an infinitive expansion (including Taylor expansion) can fail for the following reasons: not monotonic, not stable, not positive, not finite, and not reliable.
3.3.1 Convergent
@@
According to Formula (3.3.1), the condition for Formula (2.30) for $\log(x\pm\delta x)$ to converge is $P(x)\leq 1/\sigma=1/5$, which is confirmed numerically by $\log(x|pm\delta x)$ as $P(x)\lesssim 0.20088$. Beyond the upper bound $P(x)$, the expansion is not monotonic. Variance arithmetic rejects the distributional zero of $\log(0)$ inside the range of $[x-\delta x,x+\delta x]$ statistically due to the divergence of Formula (2.30) mathematically, with $\zeta(2n)$ as the connection of these two worlds.
According to Formula (3.3.1), except when c is a nature number, Formula (2.33) for $(x\pm\delta x)^{c}$ converges near $P(x)\simeq 1/\sigma=1/5$, with the upper bound $P(x)$ increasing with $c$. This is confirmed approximately by Figure 3.3. Beyond the upper bound $P(x)$, the expansion is not monotonic. Qualitatively, $\delta^{2}1/x$ converges slower than $\delta^{2}\sqrt{x}$ according to Formula (3.10) and (3.8).
$\delta x\lesssim 19.865$ for $e^{x\pm\delta x}$ regardless of $x$, while $\delta\log(x\pm\delta x)\lesssim 0.213$ regardless of $x$. Both are expected from the relation $\delta x\rightarrow P(e^{x})$ and $P(x)\rightarrow\delta\log(x)$, respectively, as shown by Formula (2.27) and (2.30).
@@
3.3.2 Stable
Conventional condition to terminate an infinitive series numerically is when the residual of an expansion is less than the unit in the last place of the result [13]. In practice, the last expansion term usually proxies for the expansion residual if either the residual is known to be less than the last expansion term, or if the residual is difficult to calculate [13]. This method no longer works in variance arithmetic because the expansion is limited to $4422n<$ 442, and because the expansion result is no longer precise. The probability of an imprecie value $x\pm\delta x$ not equal 00 is indicated by its z-statistic $\frac{|x|}{\delta x}$ [5]. If this probability is the bounding leakage $\epsilon$, Formula (3.13) gives the corresponding upper bound $\tau$ for this z-statistic. Because variance arithmetic has already guaranteed that the expansion is absolutely and monotonically decreasing, if the absolute value of the last expansion term is more than $\tau$-fold of the result uncertainty, the result is invalidated as not stable.
@@ 3.3.2 Stable
When the Taylor expansion is calculated without enough expansion terms, e.g., if $(1\pm 0.193)^{-2}$ is calculated with only $60$-term expansion, the result will be not stable.
3.3.3 Valid
@@
3.3.4 Finite
If the result of Taylor expansion is not finite, the expansion is invalidated as not finite. This usually happens when the expansion diverges so that it is usually not the reason to bound $\delta x$. For example $(1\pm 1)^{-2}$ has infinitive variance.
3.3.5 Reliable
@@ 3.3.6 Continuous
When $c$ is a nature number $n$, Formula (2.33) has $n$ terms, so that $\delta x$ has no upper bound for $(x\pm\delta x)^{n}$ unless for not finite, such as Formula (3.6) for $(x\pm\delta x)^{2}$. The result mean, variance and histogram of $(x\pm\delta x)^{c}$ are continuous around $c=n$. Table 3.3.6 shows that the result of $(1\pm\delta x)^{n\pm d}$ is very close to that of $(1\pm\delta x)^{n}$, even the former had upper bound for $\delta x$, while the latter has not.
@@
$n\pm d$ $0\pm 10^{-6}$ $1\pm 10^{-6}$ $2\pm 10^{-6}$ $3\pm 10^{-6}$
Upper Bound $\delta x$ $0.2006$ $0.2014$ $0.2018$ $0.2020$
[…]
table The result value and uncertainty of $(1\pm\delta x)^{n\pm d}$ vs $(1\pm\delta x)^{n}$, in which $n$ is a natural number, $0<d\ll 1$, and $\delta x$ is the upper bound for $(1\pm\delta x)^{n\pm d}$.
@@ 10.1 Summary of Statistical Taylor Expansion
Statistical Taylor Expansion suggests that a calculation result such as its convergence depends on input bounding factors, which may be another measurement of data quality.